OpenStudy (anonymous):

Solve for x: -4(x - 3) = 16?

mathslover (mathslover):

\[\Huge{-4(x)-4(-3)=16}\] \[\huge{-4x+12=16}\] \[\huge{-4x=4}\] \[\Huge{x=-1}\]

OpenStudy (anonymous):

Thank You MathsLover! You are really helping Me out a lot C:

mathslover (mathslover):

:) my pleasure @HelloBriBri

OpenStudy (anonymous):

I have other questions too that you could help me on if that wouldn't be a big deal?

mathslover (mathslover):

ok

OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 7x - 14 = 6x + 2 7 - 2x = 5x + 21 Answer The solution to equation #1 is the same as the solution to equation #2. The solution to equation #1 is smaller than the solution to equation #2. None of the statements above describe solutions to equations shown. The solution to equation #1 is larger than the solution to equation #2

mathslover (mathslover):

type the question as like this

mathslover (mathslover):

i mean make it another question

OpenStudy (anonymous):

Oh do you not know that one?

OpenStudy (anonymous):

Solve for y: y + 4 = 6

mathslover (mathslover):

y=2

OpenStudy (anonymous):

My Options for that are y = 3 y = -7 y = 7 y = -3

OpenStudy (anonymous):

I don't have 2 as an option :(